a glass prism has mue=1.5 and the refracting angle is 90degrees.if a ray falls on it at angle of incident of 30degrees then the angle of emergence is

Grade:9

1 Answers

Aziz Alam IIT Roorkee
9 years ago

The refractive index is n=1.5 (you have written 105, but I think so its 1.5.Let me know if I wrong.Either way you can figure if it is 105)
critical angle is ic=sin-1.667=41.80
If the incident angle is greater than this then there is total internal reflection.
Given , angle of incidence is 300
sin(i)/sin(r)=1/1.5
then sin(r)=1.5*0.5=.75
angle of refraction(r)=48.590